# Sign Spotting Disambiguation using Large Language Models

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025conference: ; ; ccs: Computing methodologies Computer vision
@@ 1. Introduction
Figure 1. System Overview. (a) Traditional sign spotters classify segments independently, relying solely on the visual encoder, which can occasionally fail. (b) Our proposed approach integrates context-aware disambiguation, leveraging preceding segments for linguistically coherent gloss sequences.
Sign languages are complex visual languages expressed through coordinated hand gestures, handshapes, facial expressions, and body posture, governed by distinct grammatical and syntactic rules (Sense, [n. d.]; National Deaf Children’s Society, 2024). With over 70 million Deaf individuals worldwide using more than 200 distinct sign languages (Núñez-Marcos et al., 2023), the need for computational models that can understand and process sign language at scale is both urgent and impactful. However, the multimodal and multi-channel nature of signing presents significant challenges for modeling, particularly when attempting to relate it to spoken language. Unlike spoken languages, sign languages are not linearly structured; they often rely on simultaneous visual cues, and exhibit fundamentally different linguistic properties, making alignment with spoken counterparts non-trivial.
Among the most widely studied tasks in the field are sign language recognition and translation, due to their direct utility in accessibility and communication. Prior work has shown that model performance on these tasks improves substantially when trained on gloss-level annotations (Chen et al., 2022b; Zhang et al., 2023; Hu et al., 2024; Chen et al., 2022a) or with the aid of large-scale sign datasets (Li et al., 2025). Glosses are an intermediate spoken language representation and are aligned at the sign level within continuous sign video; however, gloss annotation is expensive, time-consuming, and requires domain expertise. For instance, in How2Sign, annotating just 90 seconds of video with glosses took an hour (Duarte et al., 2021). Thus, existing gloss-annotated datasets are limited in scale, restricting progress in supervised methods. In contrast, large-scale datasets that pair videos with spoken language sentences offer broader coverage but lack temporal alignment and structural correspondence, as the order, grammar, and even vocabulary of spoken language differ substantially from sign language (Albanie et al., 2021; Shi et al., 2022; Tanzer and Zhang, 2024). This weak supervision introduces ambiguity and hinders translation models from effectively learning precise visual-to-linguistic mappings.
@@
Motivated by this, we propose a context-aware disambiguation framework that integrates an LLM into the sign spotting pipeline as seen in Figure 1. While our approach builds on dictionary look-up-based spotting, we enhance it by extracting top-$k$ candidate glosses for each localized segment and leveraging the LLM’s next-token prediction capabilities to evaluate gloss sequences based on linguistic coherence. We cast disambiguation as a constrained decoding task, employing beam search to explore high-probability gloss combinations informed by language priors. This enables the system to move beyond isolated sign predictions and instead produce coherent, contextually grounded gloss sequences. To further improve spotting performance, we explore a range of ensemble and fusion strategies for integrating visual cues. Overall, our approach bridges low-level visual recognition with high-level linguistic reasoning and demonstrates that incorporating LLM-based priors can significantly enhance robustness and accuracy in the presence of sign ambiguity. In summary, our main contributions are as follows:
@@ 2. Related Works
Large Language Models have marked a significant breakthrough in natural language processing, enabling remarkable performance across a wide range of linguistic tasks (Radford et al., 2018; Brown et al., 2020). State-of-the-art LLMs such as GPT (OpenAI, 2025), Claude (Anthropic, 2025), Gemma (Gosthipaty et al., 2025) and Grok (XAI, 2025) are built upon the transformer decoder architecture (Vaswani et al., 2017) and trained on massive text corpora, including Common Crawl (Vaughan, 2025), The Pile (Gao et al., 2020), and RefinedWeb (Penedo et al., 2023). Through this large-scale pretraining, LLMs acquire a broad understanding of language, making them well-suited for a variety of language understanding and generation tasks.
@@ 3.1.2. Dictionary-Based Matching
(2) $\mathbf{D}_{i}=\mathbf{F}^{I3D}_{i}\oplus\mathbf{F}^{LH}_{i}\oplus\mathbf{F}^{RH}_{i},$
@@
(3) $\text{score}(U_{x},\mathbf{D}_{i})=(1-\alpha)\cdot-\text{sim}_{\text{DTW}}(U_{x},\mathbf{D}_{i})+\alpha\cdot\text{sim}_{\text{cos}}(U_{x},\mathbf{D}_{i}),$
where $\alpha\in[0,1]$ is a hyperparameter controlling the contribution of each similarity metric. Note that DTW is a distance metric (lower values mean higher similarity); thus, we apply a negative sign to ensure higher values consistently indicate greater similarity.
@@ 3.1.3. Feature Fusion
(4) $\mathbf{S}^{\text{Late}}=\alpha\cdot\mathbf{S}^{\text{I3D}}+(1-\alpha)\cdot\mathbf{S}^{\text{RH}},$
where $\alpha\in[0,1]$ is a tunable hyperparameter that balances the influence of the I3D and RH shape features. We observed through preliminary experimentation that the inclusion of LH features provided negligible performance improvements; thus, they are excluded from this specific late fusion strategy.
@@
(5) $\mathbf{S}^{\text{Ensemble}}=\alpha\cdot\mathbf{S}^{\text{Mid}}+(1-\alpha)\cdot\mathbf{S}^{\text{I3D}},$
where $\alpha\in[0,1]$ is a tunable hyperparameter balancing the contributions of the intermediate and I3D feature modalities.
@@ 3.2.1. Gloss Candidate Generation
From the preceding dictionary look-up stage, a similarity distribution $\mathbf{S}\in\mathbb{R}^{V}$ (where $V$ is the vocabulary size) is obtained for each sign segment, indicating the relevance of each potential gloss candidate. This raw distribution undergoes normalization via a softmax function, and its Top-$k$ gloss candidates are then selected and forwarded to the LLM along with the prompt for subsequent disambiguation (Figure 5).
@@ 3.2.2. Prompt Formulation
The prompt provided to the LLM is dynamically constructed to incorporate both contextual information and candidate glosses. As seen in Figure 6, specific elements of the prompt are instantiated with variable content. For instance, the red-highlighted ”Short,” ”Secret,” and ”Make” represent the Top-$k$ gloss candidates, while the blue-highlighted ”Together we” signifies the previous sign glosses, providing contextual information.
@@
Table 1. Hyperparameter Optimization of $\alpha$ from ranges 0 to 1 for Different Fusion Equations based on Top-K Accuracy
@@ 3.2.3. Beam Search Decoding
Thus, for the $x^{\text{th}}$ sign unit, where $x\in\{1,\dots,X\}$ ($X$ denoting the total number of units), we first obtain a set of top-$k$ gloss candidates $\mathcal{C}_{x}=\{(g_{x}^{j},s_{x}^{j})\}_{j=1}^{k}$. Here, $s_{x}^{j}$, which is derived from the pre-computed dictionary similarity (e.g., $\mathbf{S}^{\text{Late}}$, $\mathbf{S}^{\text{Mid}}$, or $\mathbf{S}^{\text{Ensemble}}$), serves as the emission probability for candidate gloss $g_{x}^{j}$. Meanwhile, the LLM provides our transition probabilities $p(g_{x}|g_{1:x-1})$, indicating the likelihood of gloss $g_{x}$ given the sequence of previously selected glosses $g_{1:x-1}$, providing us with linguistic guidance.
@@
(6) $\hat{g}_{1:X}=\underset{g_{1:X}\in\prod_{x=1}^{X}\mathcal{C}_{x}}{\arg\max}\sum_{x=1}^{X}\left(\log p(g_{x}|g_{1:x-1})+\alpha s_{x}\right)$
In this objective, $\prod_{x=1}^{X}\mathcal{C}_{x}$ denotes the Cartesian product of the candidate sets. The conditional probability $p(g_{x}|g_{1:x-1})$ is derived from the frozen LLM’s logits, while the term $s_{x}$ represents the emission probability for the selected gloss $g_{x}$ at the $x^{\text{th}}$ sign unit. The hyperparameter $\alpha$ weighs the influence of the LLM’s linguistic score and the emission probability. To sufficiently explore the hypothesis space, we employ the algorithm with a beam width (BW) of $b=5$.
@@ 4.1.1. Synthetic Dictionary Construction
Existing large-scale British Sign Language (BSL) dictionaries often lack crucial, high-frequency vocabulary. To address this, we construct a synthetic BSL-based dictionary informed by word frequency research. We leverage a word frequency list derived from the Google Web Trillion Word Corpus (The Word Finder, 2024), cross-referencing it with findings from a BSL lexical frequency study (Fenlon et al., 2014), which identified the 100 most frequent BSL signs, and BSLDict (Momeni et al., 2020). We thus constructed four vocabulary lists of varying sizes: 1500, 2000, 3000, and 4373 words. The maximum size was capped at 4373, as it corresponded to the highest number of overlapping words between the Google frequency list and the BSLDict sign vocabulary. This constructed dictionary also serves as a template for potential future sign spotting applications.
@@ 4.1.2. Pseudo-Gloss Generation
To generate synthetic sign spotted units, we employ a pseudo-gloss generation strategy. We first curate a corpus of random English sentences using a random sentence generator (Tatman, 2017). Motivated by recent work in sign language translation (Wong et al., 2024), we then create pseudo-glosses by applying Part-of-Speech (POS) tagging via the SpaCy library. Only tokens tagged as “NOUN”, “NUM”, “ADV”, “PRON”, “PROPN”, “ADJ”, and “VERB” are retained, simulating the typical lexical content of sign glosses.
@@ 4.1.3. Simulating Ambiguity via Noise Injection
The top-$1$ option of the similarity distribution is replaced with another random word at a specified probability, simulating an incorrect classification by a sign spotter.
@@
From a given distribution, we randomly select $k$ semantically dissimilar words. We then artificially increase their similarity scores to a value exceeding the highest pre-existing score. This introduces significant noise, ensuring that random, irrelevant words appear in the top-$k$ candidates, thereby stringently testing the disambiguation capability of our method.
@@ 4.1.4. Evaluation and Results
We first perform HPO on the weighted sum parameters used in both the similarity computation and feature fusion stages. Specifically, we tune the balancing hyperparameter $\alpha\in[0,1]$ for four distinct combination strategies: (i) the sum of DTW and Cosine Similarity for I3D features (Eq. 3, used for $\mathbf{S}^{\text{I3D}}$ derivation), (ii) the sum of DTW and Cosine Similarity for RH features (Eq. 3, used for $\mathbf{S}^{\text{RH}}$ derivation), (iii) the late fusion strategy (Eq. 4), and (iv) the full-ensemble method (Eq. 5). Performance is measured using Top-1 and Top-5 accuracy.
Based on table 1, which summarizes the HPO results, we find that the I3D similarity computation benefits from a greater contribution from DTW. In contrast, the RH similarity computation achieves its peak Top-1 accuracy at $\alpha=0.9$, indicating a stronger reliance on cosine similarity. Regarding feature fusion, the late fusion strategy shows optimal Top-5 accuracy (0.8276) when I3D features contribute more ($\alpha=0.9$). Meanwhile, the full-ensemble method also performs strongly, with peak performances observed at $\alpha=0.6$, suggesting a slightly greater emphasis on the intermediate fusion distribution ($\mathbf{S}^{\text{Mid}}$). Overall, the explicit feature fusion methods consistently outperform the vanilla $\mathbf{S}^{\text{I3D}}$ and $\mathbf{S}^{\text{RH}}$ baselines.
@@ 4.2.1. Performance of Sign Spotters (No Disambiguation)
In contrast, approaches utilizing the dictionary-matching algorithm achieved substantially lower WERs (around 0.5). This highlights the benefits of dictionary-matching, which offers superior vocabulary flexibility and a training-free nature that allows easy incorporation of new gloss entries. Among these, the Late Fusion and Full-Ensemble methods proved most effective, achieving WERs of 0.4724 and 0.4924, respectively. Their improved performance is attributed to feature fusion techniques that enhanced discrimination for dictionary matching. Notably, the Late Fusion approach consistently performed best, aligning with its highest Top-1 Gloss Accuracy (54.51%) observed during $\alpha$ tuning (Table 1).
@@ 3. MYSELF LOVE WALKING SUNDAY AFTERNOON
Ground Truth:
WE-ALL IDEAS BOTH-FORWARD OPEN-MINDED
IN EVERYTHING WE-DO
Top-3 Prediction:
1. WE-ALL IDEAS BOTH-FORWARD OPEN-MINDED
IN EVERYTHING WE-DO
2. WE-ALL IDEAS BOTH-FORWARD OPEN-MINDED
IN EVERYTHING WE-DO
3. BOTH-WE IDEAS BOTH-FORWARD MENTAL AFFECT
EVERYTHING WE-DO
@@ 4.3.2. Qualitative Comparisons against Baseline
Baseline Spotter (Sincan et al., 2024) HEART WALL UNIVERSITY
@@
Baseline Spotter (Sincan et al., 2024) NEED SECOND WARM MOISTURE EXPLAIN EXPLAIN
@@
Baseline Spotter (Sincan et al., 2024) EXPECT MAKE GO MANY MANY
@@ 5. Conclusion
Acknowledgements.
This work was supported by the SNSF project ‘SMILE II’ (CRSII5 193686), the Innosuisse IICT Flagship (PFFS-21-47), EPSRC grant APP24554 (SignGPT-EP/Z535370/1) and through funding from Google.org via the AI for Global Goals scheme. This work reflects only the author’s views and the funders are not responsible for any use that may be made of the information it contains.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: ArXiv Version; 2025; .booktitle: Preprintdoi: xxxisbn: xxxccs: Computing methodologies Computer vision
@@ 1. Introduction
Figure 1. System Overview. (a) Traditional sign spotters classify segments independently. (b) Our proposed approach integrates context-aware disambiguation, leveraging preceding segments for linguistically coherent gloss sequences.
Sign languages are complex visual languages expressed through coordinated hand gestures, handshapes, facial expressions, and body posture, governed by distinct grammatical and syntactic rules (Sense, [n. d.]; National Deaf Children’s Society, 2024). With over 70 million Deaf individuals worldwide using more than 200 distinct sign languages (Núñez-Marcos et al., 2023), developing models to understand sign language at scale is both urgent and impactful. However, the multimodal and multi-channel nature of signing presents significant challenges for modeling, particularly when attempting to relate it to spoken language. Unlike spoken languages, sign languages are not linearly structured; they often rely on simultaneous visual cues, and exhibit fundamentally different linguistic properties, making alignment with spoken counterparts non-trivial.
Among the most widely studied tasks in the field are sign language recognition and translation, due to their direct utility in accessibility and communication (Zhou et al., 2023; Wong et al., 2024; Low et al., 2025a). Prior work has shown that model performance on these tasks improves substantially when trained on gloss-level annotations (Chen et al., 2022b; Zhang et al., 2023; Hu et al., 2024; Chen et al., 2022a; Low et al., 2025b) or with the aid of large-scale datasets (Li et al., 2025). Glosses are intermediate spoken language representations aligned at the sign level in continuous sign video; however, their annotations are costly, time-consuming, and requires expertise. For instance, annotating 90 seconds of video in How2Sign took one hour (Duarte et al., 2021). Thus, existing gloss-annotated datasets are small in scale, restricting supervised methods. In contrast, large-scale datasets, pairing videos with spoken language sentences offer broader coverage but lack temporal alignment and structural correspondence; as the order, grammar, and vocabulary of spoken language differ from sign language (Albanie et al., 2021; Shi et al., 2022; Tanzer and Zhang, 2024). This weak supervision introduces ambiguity and hinders translation models from trivially learning precise visual-to-linguistic mappings.
@@
[…]
@@ 3.1.2. Dictionary-Based Matching
(2) $\mathbf{D}_{i}=\mathbf{F}^{I3D}_{i}\oplus\mathbf{F}^{LH}_{i}\oplus\mathbf{F}^{RH}_{i}\in\mathbb{R}^{5120},$
@@
(3) $\text{score}(U_{x},\mathbf{D}_{i})=(\alpha_{\text{s}}-1)\cdot\text{sim}_{\text{DTW}}(U_{x},\mathbf{D}_{i})+\alpha_{\text{s}}\cdot\text{sim}_{\text{cos}}(U_{x},\mathbf{D}_{i}),$
where $\alpha_{\text{s}}\in[0,1]$ is a hyperparameter controlling the contribution of each similarity metric. Note that DTW is a distance metric (lower means more similarity); thus, we multiply its scalar with a negative sign to ensure higher values consistently indicate greater similarity.
@@ 3.1.3. Feature Fusion
(4) $\mathbf{S}^{\text{Late}}=\alpha_{\text{late}}\cdot\mathbf{S}^{\text{I3D}}+(1-\alpha_{\text{late}})\cdot\mathbf{S}^{\text{RH}},$
where $\alpha_{\text{late}}\in[0,1]$ is a tunable hyperparameter that balances the influence of the I3D and RH shape features. We observed through preliminary experimentation that the inclusion of LH features provided negligible performance improvements; thus, they are excluded from this specific late fusion strategy.
@@
(5) $\mathbf{S}^{\text{Ensemble}}=\alpha_{\text{ens}}\cdot\mathbf{S}^{\text{Mid}}+(1-\alpha_{\text{ens}})\cdot\mathbf{S}^{\text{I3D}},$
where $\alpha_{\text{ens}}\in[0,1]$ is a tunable hyperparameter balancing the contributions of the intermediate and I3D feature modalities.
@@ 3.2.1. Gloss Candidate Generation
[…]
@@ 3.2.2. Prompt Formulation
[…]
@@
Table 1. Hyperparameter Optimization of the $\alpha$ from ranges 0 to 1 for Different Fusion Equations based on Top-K Accuracy
@@ 3.2.3. Beam Search Decoding
[…]
@@
(6) $\hat{g}_{1:X}=\underset{g_{1:X}\in\prod_{x=1}^{X}\mathcal{C}_{x}}{\arg\max}\sum_{x=1}^{X}\left(\log p(g_{x}|g_{1:x-1})+\alpha_{\text{bs}}s_{x}\right)$
In this objective, $\prod_{x=1}^{X}\mathcal{C}_{x}$ denotes the Cartesian product of the candidate sets. The conditional probability $p(g_{x}|g_{1:x-1})$ is derived from the frozen LLM’s logits, while the term $s_{x}$ represents the emission probability for the selected gloss $g_{x}$ at the $x^{\text{th}}$ sign unit. The hyperparameter $\alpha_{\text{bs}}$ weighs the the LLM’s linguistic score and the emission probability. To sufficiently explore the hypothesis space, we employ the algorithm with a beam width (BW) of 555.
@@ 4.1.1. Synthetic Dictionary Construction
Existing large-scale British Sign Language (BSL) dictionaries often lack crucial, high-frequency vocabulary. To address this, we construct a synthetic BSL-based dictionary informed by word frequency research. We leverage a word frequency list derived from the Google Web Trillion Word Corpus (Tatman, 2017), cross-referencing it with findings from a BSL lexical frequency study (Fenlon et al., 2014), which identified the 100 most frequent BSL signs, and BSLDict (Momeni et al., 2020). We thus constructed four vocabulary lists of varying sizes: 1500, 2000, 3000, and 4373 words. The maximum size was capped at 4373, as it corresponded to the highest number of overlapping words between the Google frequency list and the BSLDict sign vocabulary. This constructed dictionary also serves as a template for potential future sign spotting applications.
@@ 4.1.2. Pseudo-Gloss Generation
To generate synthetic sign spotted units, we employ a pseudo-gloss generation strategy. We first curate a corpus of random English sentences using a random sentence generator (The Word Finder, 2024). Motivated by recent work in sign language translation (Wong et al., 2024), we then create pseudo-glosses by applying Part-of-Speech (POS) tagging via the SpaCy library. Only tokens tagged as “NOUN”, “NUM”, “ADV”, “PRON”, “PROPN”, “ADJ”, and “VERB” are retained, simulating the typical lexical content of sign glosses.
@@ 4.1.3. Simulating Ambiguity via Noise Injection
The top-111 option of the similarity distribution is replaced with another random word at a specified probability, simulating an incorrect classification by a sign spotter.
@@
[…]
@@ 4.1.4. Evaluation and Results
We first perform HPO on the weighted sum parameters used in both the similarity computation and feature fusion stages. Specifically, we tune the balancing hyperparameters $\alpha\in[0,1]$ for four distinct combination strategies: (i) the sum of DTW and Cosine Similarity for I3D features (Eq. 3, used for $\mathbf{S}^{\text{I3D}}$ derivation), (ii) the sum of DTW and Cosine Similarity for RH features (Eq. 3, used for $\mathbf{S}^{\text{RH}}$ derivation), (iii) the late fusion strategy (Eq. 4), and (iv) the full-ensemble method (Eq. 5). Performance is measured using Top-1 and Top-5 accuracy.
Based on table 1, which summarizes the HPO results, we find that the I3D similarity computation benefits from a greater contribution from DTW. In contrast, the RH similarity computation achieves its peak Top-1 accuracy at $\alpha_{\text{s}}=0.9$, indicating stronger reliance on cosine similarity. For feature fusion, the late fusion strategy shows optimal Top-5 accuracy (0.8276) when I3D features contribute more ($\alpha_{\text{late}}=0.9$). Meanwhile, the full-ensemble method also performs strongly, with peak performances observed at $\alpha_{\text{ens}}=0.6$, suggesting a slightly greater emphasis on the intermediate fusion distribution ($\mathbf{S}^{\text{Mid}}$). Overall, the explicit feature fusion methods consistently outperform the vanilla $\mathbf{S}^{\text{I3D}}$ and $\mathbf{S}^{\text{RH}}$ baselines.
@@ 4.2.1. Performance of Sign Spotters (No Disambiguation)
In contrast, approaches utilizing the dictionary-matching algorithm achieved substantially lower WERs (around 0.5). This highlights the benefits of dictionary-matching, which offers superior vocabulary flexibility and a training-free nature that allows easy incorporation of new gloss entries. Among these, the Late Fusion and Full-Ensemble methods proved most effective, achieving WERs of 0.4724 and 0.4924, respectively. Their improved performance is attributed to feature fusion techniques that enhanced discrimination for dictionary matching. Notably, the Late Fusion approach consistently performed best, aligning with its highest Top-1 Gloss Accuracy (54.51%) observed during $\alpha_{\text{late}}$ tuning (Table 1).
@@ 4.3.2. Qualitative Comparisons against Baseline
Baseline (Sincan et al., 2024) HEART WALL UNIVERSITY
@@
Baseline (Sincan et al., 2024) NEED SECOND WARM MOISTURE EXPLAIN EXPLAIN
@@
Baseline (Sincan et al., 2024) EXPECT MAKE GO MANY MANY
@@ 5. Conclusion
Acknowledgments: This work was supported by the SNSF project ‘SMILE II’ (CRSII5 193686), the Innosuisse IICT Flagship (PFFS-21-47), EPSRC grant APP24554 (SignGPT-EP/Z535370/1) and through funding from Google.org via the AI for Global Goals scheme. This work reflects only the author’s views and the funders are not responsible for any use that may be made of the information it contains.